OpenStudy (96mertcank96):

solve for w : 5/3w -1/9 =1/w

OpenStudy (anonymous):

\[\frac{ 5}{ 3w } -\frac{ 1 }{ 9 } =\frac{ 1 }{ w }\] Yes?

OpenStudy (96mertcank96):

yes